Homework 2  EE290G Intro. to MEMS  Due 2/4/97

For this homework set you will need to go to the engineering library (hopefully). Turn it in by email in concise form by midnight on 2/4/97.

1. For the following journals:

• IEEE/ASME J. Microelectromechanical Systems

• Sensors and Actuators A (Physical)

• Sensors and Materials

Answer the following questions:

(a) What are the call letters?

(b) Where are the recent issues located physically in the library?

(c) Where are the bound back issues located physically in the library?

(d) Who is the editor of the journal?

(e) Is there anyone from Berkeley on the editorial board?

2. Find another MEMS related journal and answer the same questions about it.

3. For at least 12 of the following people:

• Mark Allen

• Henry Baltes

• Bernhard Boser

• Miko Elwenspoek

• Hiro Fujita

• Toshio Fukuda

• Henry Guckel

• Roger Howe

• Greg Kovacs • Noel MacDonald • Khalil Najafi

• Kurt Petersen • Al Pisano

• Nico de Rooij

• Marty Schmidt • Steve Senturia • Isao Shimoyama • Ken Wise

Answer the following questions:

(a) Where did they go to school?

(b) Where do they work now?

4. Which was your favorite paper (or papers) that you saw while preparing this homework?

5. Where is the best place to read in the library?